The Collider Detector at Fermilab Top and Bottom Physics

Ken J. Ragan McGill University Pekka K. Sinervo University of Toronto

1. Brief description of experiment and detector 2. Physics goals and progress to date

3. IPP Participation for 1991

4. Future Prospects

The CDF Collaboration

The CDF collaboration is a group of 20 institutions

• studying p p interactions at √s = 1.8 TeV.

• Focus is on high-pT physics

Collaboration has already had a number of runs

• First engineering run with only part of the detector in 1986

• First physics run in 1987 – collected 30 pb−1 of data

– Minimum bias trigger

– 22 W → eνe candidates

– measured jet pT spectrum and inclusive particle spectrum

• Had first major run in 1988-89 – collected 4.4 pb−1 of data

– Had full L1/L2/L3 trigger

– ∼ 3000 W → eνe candidates

– 10 publications with several more in progress

• Preparing for a new run in 1991

The CDF Detector

• The CDF detector is a general purpose, solenoidal detector

• Solenoid magnet of radius ∼ 1.5 m and length 5 m generating a 1.4 T magnetic field.

• Tracking chambers inside magnetic field provide particle

momentum resolution of

\frac{σp}{p} ∼ 0.0011p

• Calorimeters surround the interaction point

• Used for electron and muon identification

• Detection of jets

• Measurement of underlying event

• Detectors outside of the calorimeter used to identify Muons

Lepton ID

Electron ID provided by several systems:

• Transverse shower profile

• Longitudinal energy deposition (had/em)

• Shower profile measurement using PWC’s imbedded in EM • Shower-track position match

• Calorimeter Energy-track momentum match

Muon ID uses drift chambers outside of hadron calorimeter

• Matching of muon track stub and track in CTC

• minimum ionizing energy deposition in calorimeter

Neutrino’s identified by presence of missing E_{T} (E̸_{T})

• Provided by almost hermetic calorimetry over |η| < 4.2

Data Collection

Focus of experiment is on high-p_{T} physics

• “interesting physics” cross section is ∼ 10 − 100 nb

• total cross section is ∼ 50 mb

• ∼2 out of every 50,000 interactions were recorded in 1988

• Decision to keep an event made by a 3-level trigger system: – L1: 90% of events are thrown out after the first 3 μs – L2: 99.9 % of remainder are thrown out after 100 μs – L3: 70% of remainder are removed using a software

algorithm

• Collected 4.4 pb^{−1} of data in 1988-89 – (5 × 10^{6} events)

Physics Results to Date

Can divide the focus of the experimental effort into

1. Electroweak Physics (W ’s and Z’s)

2. Perturbative QCD (jets)

3. Non-perturbative QCD (minimum-bias)

4. Non Standard Model Studies (SUSY, W ′’s, etc)

5. Heavy Flavour Physics (top and bottom quark studies)

Biggest problem to overcome was accessibility to data

• Solved it by making a “spin” through raw data selecting most interesting events

– Good inclusive electron or muon candidate – Good dilepton candidate

• Had access to the full data set within one month after com-

pleting the run

Will touch on a few of these areas, with primary focus the heavy flavour studies.

ElectroWeak Physics

Properties of the W and Z bosons:

• Measured the mass and width of the Z

– collaboration felt that it was important to get out an independent measurement before SLC/LEP

– used sub-sample of the best measured electrons and muons

– calibrated tracking chamber in situ and cross-calibrated

calorimeter using inclusive electron sample MZ = 91.9 ± 0.3(stat. + syst ) ± 0.2(scale ) GeV/c2

• Measured the mass and width of the W

– Difficult measurement as it uses the transverse mass

M{eν \atop T} ≡ √2ET E̸T (1 − cos ϕeν)

Use of E̸T introduces systematic uncertainties associated with entire calorimeter response and underlying event

– Reduced these by considering events with no jets

– 1130 eνe and 592 μνμ candidates MW = 79.91 ± 0.39 GeV/c2

Results in a measurement of sin2 θW = 0.231 ± 0.008

Other Electroweak studies have been completed:

• Cross section ratio

\frac{σ(W → eνe)}{σ(Z → ee)} = \frac{σ(pp → W X)}{σ(pp → Z)} \frac{Γ(W → eν)}{Γ(Z → ee)} \frac{Γ(W)}{Γ(Z)}

gives another test of the standard model

– \frac{σ(W →eνe)}{σ(Z→ee)} = 10.2 ± 0.8 ± 0.4

– Using LEP Γ(Z) measurement, can determine

Γ(W) = 2.19 ± 0.20 GeV

– If assume Nν = 3, then Mtop> 35 GeV/c2 at 95% CL.

• Z → e+ e− asymmetry provides another measure of the left and right-handed gauge couplings

• W → e+ νe asymmetry is an indirect measure of the parton distribution functions

• Direct σ(W) and σ(Z) measurements test QCD and Standard Model

It is worth noting that most of this work rests on the understanding of high-p_{T} leptons

• This work has been very collaborative involving groups from Argonne, Chicago, FNAL, Harvard, Penn and ourselves

• Efforts have been very well coordinated and analyses have been very cooperative

– Unusual in such large collaborations

There has been extensive studies of the inclusive jet data and the multijet events

• Measured inclusive jet cross section

• Compared kinematics of multi-jet events with QCD predictions (including next-to-leading order contributions).

Several groups have been searching for exotic phenomena:

• W ′’s and Z′’s

• SUSY particles

The Top Quark Search

The top quark is one of the most important physics topics to be addressed at the Tevatron

• Standard Model is riddled with inconsistencies if top quark doesn’t exist with a mass < ∼ 180 − 200 GeV/c2

• CDF has one of the best chances of discovering it in the next few years

Top Search in the 1988-89 run focused on

• Inclusive lepton final states

• Dilepton final states

• Combined multi-channel analysis

Search in Inclusive Lepton Channel

We searched for top by requiring

• A “golden” electron candidate

• At least two jets with |η| < 2.0 and ET > 10 GeV

• Require E̸T > 20 GeV to tag presence of energetic neutrino

Studied the backgrounds in detail

• Removed conversions using tracking information

• Rate of QCD fakes and interacting π+ < ∼ 10%

• Events from semi-leptonic b decays < ∼ 10%

• Primary background comes from real pp → W + jets

Removed W + jets background statistically by fitting transverse mass distribution to

\frac{dN}{dM_{T}^{eν}} = αT (M_{T}^{eν}) + βW (M_{T}^{eν}).

To verify our understanding of M_{T}^{eν}, we also studied the W + 1 jet data

• t t contribution small because of 1 jet requirement

• Good agreement confirms that we understand M{eν \atop T}

The dilepton search in the eμ channel found one candidate event

• It is a very clean event

• also has a forward μ candidate (from B?)

• not typical of primary background (Z → τ+ τ−)

• Assuming candidate is real t t, set mass limit of 28 < Mtop < 72 GeV/c2 (at 95% CL)

Subsequently extended limit to exclude up to 91 GeV/c^{2}

• Used e+ e− and μ+ μ− after removing Z’s

• Used inclusive lepton + jets by requiring a low-energy μ candidate from B decay

Bottom Quark Physics

Inclusive lepton and dilepton samples are expected to have significant contributions from b quark production

• earlier p p experiments have had difficulty isolating a clean inclusive lepton sample

– signal rates are much lower and backgrounds higher – lepton ID runs out at low momentum

• CDF has several advantages – inherent signal-to-noise is improved

– electron ID is quite robust

– backed-up by reasonably good μ system

– dilepton samples are quite clean

We have studied both the inclusive electrons and dimuon samples

Inclusive Electron Sample

The kinematics of the inclusive electron sample are what one expects from b quark semi-leptonic decay:

• Lepton ET spectrum and isolation distributions

• Missing ET distributions

• Jet activity (multiplicity and energy flow)

Most convincing argument is the observation of a D^{∘} signal around electron

• Electron charge correlated with strangeness of resulting K± b → W −cX → e−νe s X

• Rate is in good agreement with expectations

J/Psi Sample

The dimuon data sample shows a very clean J/Ψ

• Initial goal was to determine B cross section – An important “engineering” number

– A test of the next-to-leading order QCD calculations • Subsequently discovered that one could reconstruct exclusive

B decays in CDF

We start with a clean J/Ψ → μ^{+} μ^{−} signal

• Signal is already clean after the Level 2 trigger selection • QCD expectations are that:

– ∼ 60% come from B → J/Ψ

– Remainder come from χc → J/Ψ

• One goal has been to quantify this fraction

We have measured the B content using several different techniques

• Rate of tri-muon events

• Angular correlations of J/Ψ and accompanying jets

• spectrum

• Exclusive B decay modes

Exclusive Decay Modes

The most promising decay modes have been B^{±} → J/ΨK^{±}

B^{∘} → J/ΨK^{∘}

• Required K± candidates to be within 30∘ of J/Ψ

• Required high-quality tracks and good secondary vertex

Observe a signal in the J/ΨK^{±} effective mass distribution

Future Prospects for B Physics

We’ve shown we can do competitive B physics For the next run, the challenge is obtain a large sample of B’s

• Muon system is being improved – Muon threshold being reducecd

– angular coverage augmented to |η| < 1.0

– More iron being added in central region

• SVX being installed – Will give a powerful additional handle on B’s

– 1991 run will allow us to develop triggers for subsequent runs

• More effort going into di-electrons – ET threshold being lowered below 5 GeV

– inclusive trigger threshold being dropped

Hope is to increase B sample by at least an order of magnitude

Goals for 1991 CDF Run

• Detector upgrades

– Improvements to μ system

– Adding SVX and new VTX

– increasing DAQ bandwidth and L3 CPU power

• Primary physics goals

– Increase integrated luminosity by factor of > 5 – Extend top quark search to ∼ 130 − 140 GeV/c2 – Increase B sample by at least order of magnitude – Improve W and Z mass measurement

– Extend search for new physics

IPP Contributions

Commissioning of new Event Builder (EVB-II)

• production version of prototype used in 1988 run

• will have ×3 increase in data throughput – increased parallelization in readout

– improved internal control

• Assist in general DAQ commissioning and maintenance

L3 Upgrade

• We have been extensively involved in L3 effort – diagnostic system

– L3 code maintenance

• Our contribution for 1991 run will be part of the new compute engine for the L3 farm

– Currently plan on using ∼ 4 SGI 4D/280’s – IPP contribution would be one compute server

Offline software is area of third major committment

• Assist in port of offline code to UNIX

• Responsible for CDF simulation software – requires signficant mods for 1991 detector

– tuning of calorimeter response

Summary of manpower effort: Major Budget Items

Salaries  168,500
Workstations  81,500
Travel  47,400
Miscellaneous Operating  8,000
Total  305,400
L3 SGI 4D/280  143,000

Future Plans

Following Collider run scheduled for early 1994

• Expect to get another factor of ∼ 5 increase in ∫ L dt

• Major upgrade to detector – New plug/forward calorimeters (scintillator/tile)

– New DAQ system to handle 400 ns bunch crossings – Upgraded forward muon system

– Upgrade to offline data processing capability

Physics should still be compelling

• Extend top search or study it!

• Can start talking about studying CP violation in B system • Study di-vector boson production

• B-mixing studies

Provide an active physics program for mid-90’s

• Provides natural transition from Tevatron to multi-TeV physics • Will provide “test-bed” for SSC detector technologies

• CDF will continue to be competitive

Current Status of CDF Effort

CDF management is concerned about prior Canadian reaction to a CDF affiliation

• wish to see us continue as collaborators

• encouraging us to become committed to 1993 run

• requiring a Canadian committment to the CDF program be-

fore accepting us as offical collaborators

Part of this committment has already been made:

• Ken and I are clearly committed to the CDF effort

• actively recruiting additional collaborators, postdocs and students

• are fulfilling committments for 1991 run

Effort requires sufficient operating support and significant hardware committment to CDF